# Color Chiral Cherenkov radiation and energy loss in quark-gluon plasma

### v1

@@ I Introduction
The equations of motion derived from the Lagrangian (1) depend only on the gradient $\partial_{\mu}\theta$. We adopt a model of the quark-gluon plasma with spatially homogeneous and slowly time dependent $\theta$: $c_{A}\partial_{\mu}\theta=b_{0}\delta_{\mu 0}$, where $b_{0}$ is the (constant) chiral magnetic conductivity. As a result the chiral magnetic current ${\bm{j}}^{a}=b_{0}{\bm{B}}^{a}$ emerges as a source of the color magnetic field in Amper’s law.
@@ III The chiral Cherenkov radiation rate
$e_{k\lambda}=\frac{1}{\sqrt{2}}\left(0,1,\lambda^{\prime}i,\frac{k_{\perp}}{E^{\prime}}\right)\,,$ (17c)
The integral over the ${\bm{p}}^{\prime}$ in (11) is trivial considering the delta function expressing the momentum conservation. The remaining delta function in $q\to qg$ rate reads:
@@
Clearly, the argument of this delta-function can vanish only if $\lambda b_{0}>0$. We assumed in Introduction that $b_{0}>0$. Therefore, only the right-polarized gluons $\lambda>0$ can be radiated by the incident quark. Moreover, the energy conservation, expressed by the argument of the delta function (III), can be satisfied only if
@@
The delta function in $g\to gg$ rate reads:

### v2

@@ I Introduction
The equations of motion derived from the Lagrangian (1) depend only on the gradient $\partial_{\mu}\theta$. We adopt a model of the quark-gluon plasma with spatially homogeneous and slowly time dependent $\theta$: $c_{A}\partial_{\mu}\theta=b_{0}\delta_{\mu 0}$, where $b_{0}$ is the (constant) chiral magnetic conductivity. As a result the chiral magnetic current ${\bm{j}}^{a}=b_{0}{\bm{B}}^{a}$ emerges as a source of the color magnetic field in Amper’s law. That $\theta$ is a slow function of time is indicated by the parametrically large sphaleron transition time $1/(g^{4}T)$ at a given plasma temperature $T$ [17, 18, 19].
@@ III The chiral Cherenkov radiation rate
$e_{p^{\prime}\lambda^{\prime}}=\frac{1}{\sqrt{2}}\left(0,1,\lambda^{\prime}i,\frac{k_{\perp}}{E^{\prime}}\right)\,,$ (17c)
The integral over the ${\bm{p}}^{\prime}$ in (11) is trivial considering the delta-function expressing the momentum conservation. The remaining delta-function in $q\to qg$ rate reads:
@@
Clearly, the argument of this delta-function can vanish only if $\lambda b_{0}>0$. We assumed in Introduction that $b_{0}>0$. Therefore, only the right-polarized gluons $\lambda>0$ can be radiated by the incident quark. Moreover, the energy conservation, expressed by the argument of the delta-function (III), can be satisfied only if
@@
The delta-function in $g\to gg$ rate reads:
@@ V Summary and outlook
Acknowledgements.
This work was supported in part by the U.S. Department of Energy Grants No. DE-SC0023692.